OpenStudy (anonymous):

-3-√-48÷36 OPERATIONS INVOLVING SQUARE ROOTS OF NEG NUMBERS

OpenStudy (anonymous):

does this mean \[\frac{-3-\sqrt{-48}}{36}\]

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

\[\sqrt{-48}=\sqrt{16\times 3\times -1}=4\sqrt{3}\sqrt{-1}=4\sqrt{3}i\]

OpenStudy (anonymous):

so you have \[\frac{-3-4\sqrt{3}i}{36}\]

OpenStudy (anonymous):

can write in standard form as \[-\frac{1}{12}-\frac{\sqrt{3}}{9}i\]